OpenStudy (anonymous):

Write the partial fraction decomposition of the rational expression. Check your result algebraically. 5x divided by (x-4)^2 and put it step by step how did you get it.

OpenStudy (anonymous):

\[\frac{5x}{(x-4)^2}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\frac{5x}{(x-4)^2}=\frac{A}{x-4}+\frac{B}{(x-4)^2}\] is step one

OpenStudy (anonymous):

i meant add on the right

OpenStudy (anonymous):

i guess now you have to multiply that mess out

OpenStudy (anonymous):

what does A and B stand for? I have not learn this yet.

OpenStudy (anonymous):

ooooh big mistake, ignore me lets try again

OpenStudy (anonymous):

oh okay haha

OpenStudy (anonymous):

ok do you know what the goal here is? what you answer is supposed to look like?

OpenStudy (anonymous):

Uh..sadly no. My teacher have not teach me this. I believe.

OpenStudy (anonymous):

you job is to make the expression on the left look like the expression on the right i.e. you goal is to find A and B that make this work \[\frac{5x}{(x-4)^2}=\frac{A}{x-4}+\frac{B}{(x-4)^2}\]

OpenStudy (anonymous):

A and B are some numbers, some constants, and your job it to find them

OpenStudy (anonymous):

Oh, okay.

OpenStudy (anonymous):

So, how do you find it?

OpenStudy (anonymous):

if you add up on the right, you get a numerator of \[A(x-4)+B\]

OpenStudy (anonymous):

in other words, if you add on the right you get \[\frac{5x}{(x-4)^2}=\frac{A(x-4)+B}{(x-4)^2}\] this tells you that \[A(x-4)+B=5x\]

OpenStudy (anonymous):

multiply and see that \[Ax-4A+B=5x\] which means that \(A=5\) and since \(A=5\) and \(-4A+B=0\) you get \(B=20\)

OpenStudy (anonymous):

some of these steps might be confusing, if so let me know but the final answer is \[\frac{5x}{(x-4)^2}=\frac{5}{x-4}+\frac{20}{(x-4)^2}\]

OpenStudy (anonymous):

Thank you so much! I really understand it now!

OpenStudy (anonymous):

yw